# Influence of Non-extensivity on the drag and diffusion coefficients of hadronic matter

### v1

@@ Abstract
In this work, we investigate the drag and diffusion coefficients of various hadrons propagating through a hadronic thermal bath by employing the Fokker–Planck equation within the framework of Tsallis non-extensive statistics. The non-extensive parameter $q$ accounts for the deviation from equilibrium and provides a more realistic description of the medium that is not perfectly thermalized. The hadronic bath, consisting of various mesonic and baryonic species, is characterized by different mass cutoffs that control the spectral composition of the medium. Our analysis shows that both the drag $F$ and momentum diffusion coefficients $\Gamma$ increases exponentially with temperature and increases systematically with increasing $q$ and mass cutoff. The spatial diffusion coefficient $D_{x}$ exhibits a decreasing trend with temperature $T$, $q$ and mass cutoff which highlights the significant influence of non-equilibrium effects and hadronic composition on the transport behaviour of hadrons, offering valuable insights into the thermal and dynamical properties of the hadronic phase preceding freeze-out in heavy-ion collisions. Additionally, we have studied the relaxation time of heavy mesons such as $D_{0}$, $J/\psi$ and $\Upsilon$. We found that the heavier mesons relaxed later in comparison to the lighter mesons in the hadronic medium.
@@ I Introduction
Relativistic heavy-ion collision experiments at major facilities such as RHIC and the LHC offer an exceptional opportunity to explore the properties of strongly interacting matter at extreme temperatures and densities. In these high-energy environments, ordinary nuclear matter undergoes a remarkable transformation from a confined hadronic phase to a deconfined state composed of quarks and gluons, known as the quark–gluon plasma (QGP) Niida and Miake (2021); Iancu (2014); Steinberg (2011); Wyslouch (2011). This transition marks a fundamental change in the behaviour of matter governed by Quantum Chromodynamics (QCD). Following the collision, the initially formed hot and dense QGP rapidly expands and cools, eventually hadronizing into a system dominated by interacting hadrons. During this hadronic phase, the medium remains strongly coupled, with significant interactions among its constituents before the system reaches thermal and chemical freeze-out He et al. (2012); Cao et al. (2015); Becattini et al. (2014). A comprehensive understanding of this evolution—from the early deconfined stage to the late hadronic phase—requires detailed knowledge of the transport properties such as drag, diffusion and viscosity. These microscopic quantities play a crucial role in determining the rate of thermalization, collective flow and energy loss mechanisms within the medium. By studying these properties, researchers can better characterize the dynamics of the QGP, the nature of the phase transition, and the emergent behaviour of strongly interacting matter created in relativistic heavy-ion collisions.
@@
Heavy mesons are excellent probes for investigating the quark–gluon plasma (QGP) due to their distinct production and interaction characteristics. Heavy quarks, such as charm and bottom are produced in the initial hard scatterings of these collisions, well before the QGP is formed, allowing them to experience the entire evolution of the medium — from the deconfined QGP phase through hadronization to the hadronic stage. Their large masses suppress thermal production, ensuring that any observed modifications in their spectra or flow patterns arise solely from interactions with the medium rather than from new production processes.Cao et al. (2015); He et al. (2012); Svetitsky (1988) The propagation of heavy quarks through the QGP involves numerous soft scatterings with medium constituents, leading to momentum degradation (drag) and momentum broadening (diffusion), which can be effectively described by the Fokker–Planck transport approach. These transport coefficients directly encode the strength of the coupling between heavy quarks and the QGP. Furthermore, the resonant interactions near the critical temperature $T_{c}$ enhance the coupling between heavy quarks and the medium, resulting in stronger drag and diffusion. During hadronization, heavy quarks recombine with light quarks through coalescence processes that retain imprints of the medium’s dynamics. Consequently, the study of heavy-meson transport and diffusion coefficients provides a sensitive means of characterizing the coupling strength, microscopic dynamics, and transport properties of both the quark–gluon plasma and the hadronic matter phases.
@@
Recent studies have utilized a range of QCD-inspired theoretical approaches to investigate the temperature dependence of drag and diffusion coefficients for hadrons and heavy quarks. In reference Laine (2011), Heavy Meson Chiral Perturbation Theory was employed to analyze diffusion coefficient, while reference He et al. (2011) examined transport coefficients using empirical elastic scattering amplitudes of $D$ mesons with thermal hadrons. Additionally, effective field theory–based calculations of drag and diffusion were carried out in references Ghosh et al. (2011); van Hees and Rapp (2005) and reference He et al. (2011) utilized vacuum scattering amplitudes derived from effective Lagrangians. More recently, transport properties have also been explored within the Color String Percolation Model (CSPM) Goswami et al. (2023b) and the Van der Waals Hadron Resonance Gas (VDWHRG) framework Goswami et al. (2023a). Traditionally, these studies are carried out assuming a medium in perfect thermal equilibrium, described by Boltzmann–Gibbs statistics. However, experimental observations indicate that the particle momentum spectra in heavy-ion collisions often exhibit power-law tails, suggesting significant non-equilibrium effects. To capture such deviations, the Tsallis non-extensive statistics has emerged as a powerful generalization of conventional thermodynamicsAkhil and Tiwari (2024); Singh et al. (2024). The non-extensive parameter $q$ quantifies the degree of deviation from equilibrium, with $q=1$ recovering the Boltzmann–Gibbs limit and $q>1$ representing systems with long-range correlations and memory effectsTsallis (1988); Kyan and Monnai (2022); Zheng et al. (2015); Singh et al. (2024). In this framework, the non-extensive parameter $q$ modifies the particle distribution functions, directly affecting the drag and diffusion coefficients. Hence, their study under Tsallis statistics reveals key insights into the non-equilibrium transport behaviour of hadrons in a thermal hadronic medium.
@@ II.1 The Drag and Diffusion of Hadron gas
To investigate the transport dynamics of hadrons in QGP produced in relativistic heavy-ion collisions, we employ the Fokker–Planck equation as a theoretical framework which provides a microscopic description of the stochastic evolution of the hadrons momentum distribution due to its interaction with the thermalized medium constituents which is light quarks. The dissipative and fluctuating aspects of the medium-induced transport which is explained by drag and diffusion coefficients, play a central role in quantifying the degree of thermalization and energy loss mechanisms of hadrons within the QGP.
The Fokker-Planck equation in three dimension is given by Abreu et al. (2011)
@@
The standard diffusion equation is given by the equation
@@ III Results and Discussions
In this section, we present a detailed analysis of the drag and diffusion coefficients of hadrons propagating through a hadronic thermal medium. The transport coefficients have been evaluated using the Fokker–Planck formalism within the framework of non-extensive Tsallis statistics, which provides a generalized description of systems that deviate from ideal equilibrium conditions. The non-extensive parameter $q$ quantifies the degree of deviation from equilibrium and serves as a crucial probe of the medium’s thermodynamic state, offering a more realistic representation of the hadronic environment formed in relativistic heavy-ion collisions where complete thermalization is rarely achieved. The interactions of hadrons with the surrounding light and heavy quarks constitute the effective thermal bath, whose microscopic properties are influenced by the inclusion of higher resonant statesvan Hees and Rapp (2005). To account for these contributions, different hadronic mass cutoffs have been implemented, allowing a systematic study of how heavier resonances modify the thermal density, scattering rates and overall transport behaviour of the medium. The mass cutoff thus plays a significant role in determining the effective strength of interactions between the probe particle and the medium constituentsCastorina et al. (2010). The variation of the drag and diffusion coefficients with the Tsallis parameter $q$ reveals the sensitivity of the transport properties to non-equilibrium effects in the medium. Thus, the Tsallis non-extensive parameter $q$ and the chosen mass cutoff provides important insights into the interplay between thermal and non-equilibrium effects in the strongly interacting hadronic phase.
@@
The drag coefficient exhibits an exponential rise with increasing temperature, reflecting the growing strength of interactions between the propagating hadron and the constituents of the hadronic medium. As the temperature increases, the thermal population and average momentum of hadrons become significantly higher, leading to an enhanced rate of collisions and stronger momentum dissipation. This behaviour signifies that the medium becomes increasingly dense and interactive as it approaches the critical temperature Bazavov and others (2014), indicating the onset of strong coupling in the hadronic phase.
@@
The drag coefficient also increases with the hadronic mass cutoff as seen in fig 1(b), which determines the extent to which heavier resonant states are included in the thermal spectrum. A higher cutoff incorporates additional massive resonances, thereby increasing the effective density and the number of available scattering channels in the medium. The inclusion of these heavy states enhances the overall interaction strength between the heavy resonant states and the surrounding hadrons, leading to a noticeable rise in the drag coefficient. This trend underscores the role of hadronic resonances in strengthening transport interactions and modifying the microscopic dynamics of the thermal bath near critical temperature Bazavov and others (2014).
@@
The relaxation time $\tau$ = $F^{-1}$ of heavy mesons is presented in figure 3 together with their corresponding ratios. A clear increase in the relaxation time is observed with increasing mesons mass, indicating that heavier mesons require a longer time to relax in the hadronic medium. Consequently, the heavier mesons are unlikely to fully relax before exiting the hadronic medium. Furthermore, the ratios of relaxation time remains nearly constant over the considered temperature range. Quantitatively, the relaxation-time ratio lies in the range 2.3-2.9 for $J/\psi$ relative to $D_{0}$ and a ratio of 1.6-1.9 is obtained for $\Upsilon$ relative to $D_{0}$.
Figure 4 illustrates the variation of the momentum diffusion coefficient of the $D_{0}$ meson with temperature for different values of the Tsallis parameter $q$ and hadronic mass cutoffs using Eq.(20). We notice an exponential increase in momentum diffusion along with temperature align with Ghosh et al. (2011); Torres-Rincon et al. (2022) and a systematic increase with parameter $q$ and mass cutoffs. The coefficient exhibits a pronounced exponential rise with increasing temperature, reflecting the enhanced random momentum transfer between mesons and the surrounding hadrons at higher thermal energies. As the medium becomes hotter and denser, the frequency and strength of particle collisions increases, amplifying stochastic interactions and momentum fluctuations. This behaviour indicates a transition towards a more strongly coupled regime near the critical temperature, where thermal agitation and scattering effect dominates the transport dynamics. The momentum diffusion coefficient also increases systematically with the Tsallis non-extensive parameter $q$ can be seen in fig 4(a). A higher value of $q$ represents stronger deviation from equilibrium and a broader momentum distribution within the medium. This extended high-energy tail enhances the probability of large-momentum transfer processes, thereby intensifying the random kicks imparted to the propagating mesons. A similar increasing trend is observed with the hadronic mass cutoff in fig 4(b) which defines the upper limit of resonant states contributing to the thermal bath. As the mass cutoff increases, more massive hadronic resonances are included, enriching the medium’s particle content and interaction strength. The presence of the heavier states provides additional scattering channels and enhances the overall collision rate, leading to greater momentum diffusion. This dependence highlights the significant role of the hadronic spectrum in governing the transport behaviour of hadrons and further supports the picture of a dense and strongly interacting medium.
@@ IV Summary and Outlook
    The drag coefficient exhibits an exponential rise with temperature, suggesting that the interaction between the probe particle and the medium becomes increasingly stronger at higher temperatures. This behaviour reflects the enhanced coupling strength of the system. Such increase signifies more frequent and intense scattering processes within the quark-gluon plasma.
@@
    In contrast to the drag and momentum diffusion coefficients, the spatial diffusion coefficient decreases with increasing temperature, non-extensive parameter $q$, and hadronic mass cutoff. This decline indicates that the particle’s mobility within the medium becomes progressively limited under such conditions. The stronger interactions at higher temperature and with heavier resonances lead to reduced spatial transport. Consequently, the meson experiences enhanced confinement and a higher degree of localization within the quark-gluon medium.
@@
  7. 7.
    These overall trend point towards the formation of a dense and strongly interacting hadronic medium. The enhanced coupling and reduced particle mobility reflects the onset of strong correlations and collective behaviour within the medium.

### v2

@@ Abstract
In this work, we investigate the drag and diffusion coefficients of various hadrons propagating through a hadronic thermal bath by employing the Fokker–Planck equation within the framework of Tsallis non-extensive statistics. The non-extensive parameter $q$ accounts for the deviation from equilibrium and provides a more realistic description of the medium that is not perfectly thermalized. The hadronic bath, consisting of various mesonic and baryonic species, is characterized by different mass cutoffs that control the spectral composition of the medium. Our analysis shows that both the drag $F$ and momentum diffusion coefficients $\Gamma$ increases exponentially with temperature and increases systematically with increasing $q$ and mass cutoff. The spatial diffusion coefficient $D_{x}$ exhibits a decreasing trend with temperature $T$, $q$ and mass cutoff which highlights the significant influence of non-equilibrium effects and hadronic composition on the transport behaviour of hadrons, offering valuable insights into the thermal and dynamical properties of the hadronic phase preceding freeze-out in heavy-ion collisions. Additionally, we have studied the relaxation time of heavy mesons such as $D_{0}$, $J/\psi$ and $\Upsilon$. We found that the heavier mesons relaxed later in comparison to the lighter mesons in the hadronic medium at a particular temperature.
@@ I Introduction
Relativistic heavy-ion collision experiments at major facilities such as RHIC and the LHC offer an exceptional opportunity to explore the properties of strongly interacting matter at extreme temperatures and densities. In these high-energy environments, ordinary nuclear matter undergoes a remarkable transformation from a confined hadronic phase to a deconfined state composed of quarks and gluons, known as the quark–gluon plasma (QGP) Niida and Miake (2021); Iancu (2014); Steinberg (2011); Wyslouch (2011). This transition marks a fundamental change in the behaviour of matter governed by Quantum Chromodynamics (QCD). Following the collision, the initially formed hot and dense QGP rapidly expands and cools, eventually hadronizing into a system dominated by interacting hadrons. During this hadronic phase, the medium remains strongly coupled, with significant interactions among its constituents before the system reaches thermal and chemical freeze-out He et al. (2012); Cao et al. (2015); Becattini et al. (2014). A comprehensive understanding of this evolution from the early deconfined stage to the late hadronic phase requires detailed knowledge of the transport properties such as drag, diffusion and viscosity. These microscopic quantities play a crucial role in determining the rate of thermalization, collective flow and energy loss mechanisms within the medium. By studying these properties, researchers can better characterize the dynamics of the medium, the nature of the phase transition, and the emergent behaviour of strongly interacting matter created in relativistic heavy-ion collisions.
@@
Heavy mesons are excellent probes for investigating the strongly interacting matter due to their distinct production and interaction characteristics. Heavy quarks, such as charm and bottom are produced in the initial hard scatterings of these collisions, well before the QGP is formed, allowing them to experience the entire evolution of the medium from the deconfined QGP phase through hadronization to the hadronic stage. Their large masses suppress thermal production, ensuring that any observed modifications in their spectra or flow patterns arise solely from interactions with the medium rather than from new production processes.Cao et al. (2015); He et al. (2012); Svetitsky (1988) The propagation of heavy mesons through the hadronic medium involves numerous soft scatterings with medium constituents, leading to momentum degradation (drag) and momentum broadening (diffusion), which can be effectively described by the Fokker–Planck transport approach. These transport coefficients directly encode the strength of the coupling between heavy mesons and the hadronic medium. Consequently, the study of heavy mesons transport and diffusion coefficients provides a sensitive means of characterizing the coupling strength, microscopic dynamics, and transport properties of the hadronic matter.
@@
Recent studies have utilized a range of QCD-inspired theoretical approaches to investigate the temperature dependence of drag and diffusion coefficients for hadrons and heavy mesons. In reference Laine (2011), Heavy Meson Chiral Perturbation Theory was employed to analyze diffusion coefficient, while reference He et al. (2011) examined transport coefficients using empirical elastic scattering amplitudes of $D$ mesons with thermal hadrons. Additionally, effective field theory–based calculations of drag and diffusion were carried out in references Ghosh et al. (2011); van Hees and Rapp (2005). In reference He et al. (2011) vacuum scattering amplitudes derived from effective Lagrangians are utilized to study the drag and diffusion coefficients. More recently, transport properties have also been explored within the Color String Percolation Model (CSPM) Goswami et al. (2023b) and the Van der Waals Hadron Resonance Gas (VDWHRG) framework Goswami et al. (2023a). Traditionally, these studies are carried out assuming a medium in perfect thermal equilibrium, described by Boltzmann–Gibbs statistics. However, experimental observations indicate that the particle momentum spectra in heavy-ion collisions often exhibit power-law tails, suggesting significant non-equilibrium effects. To capture such deviations, the Tsallis non-extensive statistics has emerged as a powerful generalization of conventional thermodynamicsAkhil and Tiwari (2024); Singh et al. (2024). The non-extensive parameter $q$ quantifies the degree of deviation from equilibrium, with $q=1$ recovering the Boltzmann–Gibbs limit and $q>1$ representing systems with long-range correlations and memory effectsTsallis (1988); Kyan and Monnai (2022); Zheng et al. (2015); Singh et al. (2024). In this framework, the non-extensive parameter $q$ modifies the particle distribution functions, directly affecting the drag and diffusion coefficients. Hence, the study under Tsallis statistics reveals key insights into the non-equilibrium transport behaviour of hadrons in a thermal hadronic medium.
@@ II.1 The Drag and Diffusion of Hadron gas
To investigate the transport dynamics of hadrons produced in relativistic heavy-ion collisions, we employ the Fokker–Planck equation as a theoretical framework which provides a microscopic description of the stochastic evolution of the hadrons momentum distribution due to its interaction with the thermalized medium constituents. The dissipative and fluctuating aspects of the medium induced transport is explained by drag and diffusion coefficients which play a central role in quantifying the degree of thermalization and energy loss mechanisms of hadrons within the medium. The Fokker-Planck equation in three dimension is given by Abreu et al. (2011),
@@
The standard diffusion equation is given as,
@@ III Results and Discussions
In this section, we present a detailed analysis of the drag and diffusion coefficients of hadrons propagating through a hadronic thermal medium. The transport coefficients have been evaluated using the Fokker–Planck formalism within the framework of non-extensive Tsallis statistics, which provides a generalized description of systems that deviate from ideal equilibrium conditions. The non-extensive parameter $q$ quantifies the degree of deviation from equilibrium and serves as a crucial probe of the medium’s thermodynamic state, offering a more realistic representation of the hadronic environment formed in relativistic heavy-ion collisions where complete thermalization is rarely achieved. The interactions of hadrons with the surrounding medium which constitute the effective thermal bath, whose microscopic properties are influenced by the inclusion of higher resonant statesvan Hees and Rapp (2005). To account for these contributions, different hadronic mass cutoffs have been implemented, allowing a systematic study of how heavier resonances modify the thermal density, scattering rates and overall transport behaviour of the medium. The mass cutoff thus plays a significant role in determining the effective strength of interactions between the probe particle and the medium constituentsCastorina et al. (2010). The variation of the drag and diffusion coefficients with the Tsallis parameter $q$ reveals the sensitivity of the transport properties to non-equilibrium effects in the medium. Thus, the Tsallis non-extensive parameter $q$ and the chosen mass cutoff provides important insights into the interplay between thermal and non-equilibrium effects in the strongly interacting hadronic phase.
@@
The drag coefficient exhibits an exponential rise with increasing temperature, reflecting the growing strength of interactions between the propagating hadron and the constituents of the hadronic medium. As the temperature increases, the thermal population and average momentum of hadrons become significantly higher, leading to an enhanced rate of collisions and stronger momentum dissipation.
@@
The drag coefficient also increases with the hadronic mass cutoff as seen in fig 1(b), which determines the extent to which heavier resonant states are included in the thermal spectrum. A higher cutoff incorporates additional massive resonances, thereby increasing the effective density and the number of available scattering channels in the medium. The inclusion of these heavy states enhances the overall interaction strength between the heavy resonant states and the surrounding hadrons, leading to a noticeable rise in the drag coefficient. This trend underscores the role of hadronic resonances in strengthening transport interactions and modifying the microscopic dynamics of the thermal bath.
@@
The relaxation time $\tau$ = $F^{-1}$ of heavy mesons is presented in figure 3 together with their corresponding ratios. A clear increase in the relaxation time is observed with increasing mesons mass, indicating that heavier mesons require a longer time to relax in the hadronic medium. Consequently, the heavier mesons are unlikely to fully relax before exiting the hadronic medium. Furthermore, the ratios of relaxation time remains nearly constant over the considered temperature range. Quantitatively, the relaxation-time ratio lies in the range 2.3-2.9 for $\Upsilon$ relative to $D_{0}$ and a ratio of 1.6-1.9 is obtained for $J/\psi$ relative to $D_{0}$.
Figure 4 illustrates the variation of the momentum diffusion coefficient of the $D_{0}$ meson with temperature for different values of the Tsallis parameter $q$ and hadronic mass cutoffs using Eq.(20). We notice an exponential increase in momentum diffusion along with temperature align with Ghosh et al. (2011); Torres-Rincon et al. (2022) and a systematic increase with parameter $q$ and mass cutoffs. The coefficient exhibits a pronounced exponential rise with increasing temperature, reflecting the enhanced random momentum transfer between mesons and the surrounding hadrons at higher thermal energies. As the medium becomes hotter and denser, the frequency and strength of particle collisions increases, amplifying stochastic interactions and momentum fluctuations.
The momentum diffusion coefficient also increases systematically with the Tsallis non-extensive parameter $q$ can be seen in fig 4(a). A higher value of $q$ represents stronger deviation from equilibrium and a broader momentum distribution within the medium. This extended high-energy tail enhances the probability of large-momentum transfer processes, thereby intensifying the random kicks imparted to the propagating mesons. A similar increasing trend is observed with the hadronic mass cutoff in fig 4(b) which defines the upper limit of resonant states contributing to the thermal bath. As the mass cutoff increases, more massive hadronic resonances are included, enriching the medium’s particle content and interaction strength. The presence of the heavier states provides additional scattering channels and enhances the overall collision rate, leading to greater momentum diffusion. This dependence highlights the significant role of the hadronic spectrum in governing the transport behaviour of hadrons and further supports the picture of a dense and strongly interacting medium.
@@ IV Summary and Outlook
    The drag coefficient exhibits an exponential rise with temperature, suggesting that the interaction between the probe particle and the medium becomes increasingly stronger at higher temperatures. This behaviour reflects the enhanced coupling strength of the system. Such increase signifies more frequent and intense scattering processes within the medium.
@@
    In contrast to the drag and momentum diffusion coefficients, the spatial diffusion coefficient decreases with increasing temperature, non-extensive parameter $q$, and hadronic mass cutoff. This decline indicates that the particle’s mobility within the medium becomes progressively limited under such conditions. The stronger interactions at higher temperature and with heavier resonances lead to reduced spatial transport. Consequently, the meson experiences enhanced confinement and a higher degree of localization within the medium.